2. The Garble Corporation produces brass rings for its assembly of Fancy Garbles. The firm operates 220 days a year and uses the rings at a steady rate of 50 per day. Rings can be produced at a rate of 200 per day. Annual storage cost is $2 per ring, and the machine setup cost is $70 per run.

a) determine the economic run quantity

b) how many runs per year will there be?

c) what is the average inventory on hand?
Answer #1

Answers are in BOLD

Daily production rate, p 200
Daily demand/usage rate, d 50
operation days, o= 220
Annual demand, D= d*o 11000
cost per order, S $70.00
holding cost per year, H= $2.00
Answer a Economic run quantity, Q= SQRT(2*S*D/(H*(1-d/p))) 1013.25
length of production run in days, t t= Q/p 5
Answer b #Runs per year = Annual demand/ordered quantity, D/Q 10.86
Maximum inventory level, pt - dt Total produced during the product run- total used during the production run 759.93
Answer c Average inventory level Maximum inventory level/2 379.97